OpenStudy (anonymous):

Solve, showing all steps: 5(x – 3) – 2(2x + 4) = -2(x – 2) I tried this but ended up with a completely wrong answer...which was 3.67, but I know the answer is 9...How do I get that?

OpenStudy (jack1):

5(x – 3) – 2(2x + 4) = -2(x – 2) \[5(x – 3) – 2(2x + 4) = -2(x – 2)\] \[5x – 15 – 2(2x + 4) = -2(x – 2)\] \[5x – 15 – 4x - 8 = -2(x – 2)\] \[5x – 15 – 4x - 8 = -2x +4\]

OpenStudy (jack1):

x-23 = -2x +4 3x = 27 x = 9

OpenStudy (anonymous):

Shouldnt it be 4x + 8?

OpenStudy (jack1):

- everything in the brackets

OpenStudy (anonymous):

That's where I messed up I guess...

OpenStudy (anonymous):

Thanks a lot though

OpenStudy (jack1):

hall good dude